# LoX: Low-Rank Extrapolation Robustifies LLM Safety Against Fine-tuning

### v2

@@ 1 Introduction
Based on this insight, we propose Low-Rank Extrapolation (LoX), a simple yet effective training-free method for enhancing the robustness of safety-aligned LLMs. LoX extrapolates the subspace of the alignment weight updates to strengthen LLMs against fine-tuning (see Fig. 1). Through extensive experimentation, we demonstrate that our method can improve robustness against both benign and malicious fine-tuning attacks, without compromising the model’s ability to be further fine-tuned for new tasks. Furthermore, we present detailed ablation studies on the effects of the chosen subspace and the scaling coefficient. Interestingly, when inspecting the safety landscape, we observe that LoX moves the aligned model away from a narrow valley, where small perturbations can easily degrade safety, into a flatter and, therefore, safety-robust region. Our contributions are:
@@ 2 Related Work
Mitigations of Finetuning Attack. After the risks of fine-tuning were identified, several mitigation strategies were proposed (Huang et al., 2024c; Rosati et al., 2024; Wang et al., 2024; Yi et al., 2024), including (a) modifying the alignment stage to account for robustness against fine-tuning attacks (Rosati et al., 2024; Liu et al., 2024; Huang et al., 2024b; d); (b) modifying the fine-tuning stage to mitigate safety degradation (Wang et al., 2024; Zong et al., 2024; Bianchi et al., 2023; Mukhoti et al., 2023); and (c) recovering the safety from already fine-tuned models (Yi et al., 2024; Huang et al., 2024a). Another line of work focuses on improving robustness during the fine-tuning stage through techniques such as retaining base model features or applying adversarial training (Dong et al., 2021; Hou et al., 2022; Kim et al., 2023). Additional discussion regarding these methods is presented in Appendix A.
@@ 3.2 Low-Rank Hypothesis for Safety Model Updates
where the columns of $U$ and $V$ are the left-singular vectors and right-singular vectors, respectively. The matrix $S$ is diagonal and contains the singular values of $M$, ordered from the largest to the smallest along the main diagonal. Based on SVD, one can decompose $M$ into a summation of $r$ matrices with a rank of $1$ and consequently
@@ 3.3 Fine-tuning Vanishes Safety Subspaces
Setting. We apply direct preference optimization (DPO) to safety align LLaMA-2-7B on the HH-RLHF dataset (Bai et al., 2022). We conduct multiple experiments with 22.5k, 32.8k, 45k, and 65.6k examples using the OpenRLHF codebase (Hu et al., 2024). Subsequently, we fine-tune the aligned models on GSM8K (Cobbe et al., 2021b), a dataset curated for mathematical reasoning. See Appendix B for hyper-parameter details.
Evaluation. In addition to our proposed metrics (i.e., $R_{\text{align}}$ and $R_{\text{ft}}$), we compute the Attack Success Rate (ASR) using the first $100$ examples from AdvBench (Zou et al., 2023a). This allows us to quickly iterate on multiple models. Following the methodology described in Qi et al. (2023), we perform a GPT-based evaluation with a revised prompt (presented in Appendix H). We use multiple values of $k$ ($10,100,500$, and $2000$) to compute the projections ${\mathrm{Proj}}_{{k}}{(.)}$. The resulting values for $R_{\text{ft}}$ and $R_{\text{align}}$ are averaged across all weight matrices.
Figure 2: Comparison of ASR difference before and after fine-tuning on GSM8K (“Safety Degradation”) versus the ratio $R_{\text{ft}}/R_{\text{align}}$. Marker size represents the number of alignment examples ($22.5$k, $32.8$k, $45$k, $65.6$k). Higher $R_{\text{ft}}/R_{\text{align}}$ correlates with lower safety degradation, suggesting the role of top-ranks in safety robustness.
Results. Firstly, as shown in Fig. 2, safety robustness increases as the number of data used to align the models increases. For example, the model aligned with the fewest number of examples ($22.5$k) achieves $53\%$ ASR after fine-tuning, while the model aligned with the largest number of examples ($65.6$k) achieves only $15\%$ ASR.
Secondly, we find that in all cases $R_{\text{ft}}/R_{\text{align}}<1$ (so $R_{\text{ft}}<R_{\text{align}}$), which indicates that the proportion of safety knowledge decreases after fine-tuning. Furthermore, we observe a correlation between safety robustness and the ratio $R_{\text{ft}}/R_{\text{align}}$. For low values of $k$, large variations in $R_{\text{ft}}/R_{\text{align}}$ are correlated with larger variations in the ASR. As $k$ increases, changes in $R_{\text{ft}}/R_{\text{align}}$ become less significant, highlighting the importance of the top-ranks on the safety robustness. The exact values for the ASR, $R_{\text{ft}}$ and $R_{\text{align}}$ are presented in Appendix D.
@@ 4 New Method: Low-Rank Extrapolation
Based on the insights from Section 3, we propose a new method to enhance the safety robustness of an aligned model. To maximize the flexibility and compatibility of our method, we target a general defense without making assumptions about the fine-tuning mechanisms. This new method, called Low-rank Extrapolation (LoX), strengths the safety subspaces in LLMs by extrapolating the top-ranks of $\Delta W_{align}$. More specifically, we define $W_{up}$ as
@@
where $\theta_{r}=\{W_{\text{base}}^{i}+{\mathrm{Proj}}_{{r}}{(\Delta W^{i}_{\text{align}})}\}_{i=i}^{L}$, $\text{ASR}(\theta)$ denotes the Attack Success Rate of the model of parameters $\theta$, and $\rho\in\mathbb{R}$ is a threshold which we set to $0.01$. We highlight the properties of LoX:
@@ 5.1 LoX Robustifies LLMs Against Fine-tuning Attacks
To evaluate the efficacy of LoX, we fine-tune our model, aligned with $65.6$k examples of the HH-RLHF dataset (Bai et al., 2022), on a variety of benign and malicious fine-tuning tasks, both before and after applying LoX. We use $k=6$ (the effective rank) and set $\alpha=1.25$. The experiment for determining the effective rank is provided in Appendix C.
Attack Methods. In addition to the GSM8K fine-tuning described in Section 3.3, we also evaluate our method on Alpaca (Taori et al., 2023), Dolly (Conover et al., 2023), Identity Shifting Attack (Qi et al., 2023) and Pure Bad (Qi et al., 2023) tasks. For details about the datasets and hyper-parameters, see Appendix B.
@@
Results. As shown in Table 1, LoX improves the models’ robustness across a wide range of datasets, strengthening their safety after fine-tuning. We observe the largest reduction in ASR brought by LoX (up to $54\%$) on the Pure Bad task, followed by a reduction of $45\%$ in ASR on the Dolly fine-tuning, notably without reducing Helpfulness . On GSM8K, LoX also significantly increases the safety robustness with only minimal affect to accuracy. For instance, applying LoX to the GSM8K dataset achieves a reduction of 11% in ASR, while only a decrease of $0.6\%$ in accuracy. This underscores the efficacy of LoX in enhancing safety robustness.
Fig. 3 shows that the improvements in safety robustness are accompanied by increases in $R_{ft}$, in 5 different tasks. For example, while $R_{ft}$ increases by $0.247$, the ASR decreases by $23\%$ in the Alpaca task. Interestingly, the ASR decreases more dramatically on fine-tuning tasks that are more damaging to safety (Dolly, Identity Shifting and Pure Bad) than on the other tasks (GSM8K, Alpaca). Specifically, $R_{ft}$ changes similarly on Alpaca as on the task of Identity Shifting, Pure Bad and Dolly, but the ASR changes are more significant on the last three. By the definition of $R_{ft}$ in Section 3.2, the result suggests that increasing the safety top-rank norm leads to improved safety robustness.
@@ 5.2 Ablation Study
Evaluation with Stronger Attacks. A natural question that arises is whether the obtained safety robustness can be compromised by strengthening the attack. To examine this, we conduct additional experiments on the Dolly and GSM8K datasets, using varying learning rates and training for additional epochs. For Dolly fine-tuning, we use learning rates of $2\times 10^{-5}$ and $5\times 10^{-5}$, and for GSM8k, $5\times 10^{-5}$ and $1\times 10^{-4}$. Dolly is evaluated as in Section 5.1, while GSM8k accuracy is reported on 100 examples of the test set, for quick checkpoint assessment. Training runs for 6 epochs.
As shown in Fig. 4, LoX improves the model’s robustness across all epochs and fine-tuning settings, with the exception of the GSM8K with $10^{-4}$ learning rate. In this case, LoX performs identically to the baseline in epochs $1$ and $6$, but worse in epoch $2$. When inspecting the utility of the models, we observe very similar performance between LoX and the baseline, indicating that LoX does not negatively impact the model’s ability to be further adapted to new tasks.
When comparing the best-performing models, we observe that the baseline GSM8K model with the highest performance is achieved at epoch $3$ with a learning rate of $5\times 10^{-5}$, which has an accuracy of $34\%$ and an ASR of $16\%$. The highest accuracy LoX  GSM8K model is obtained at epoch 2, with an accuracy of $33\%$ and an ASR of $5\%$. This demonstrates an improvement of $11\%$ in robustness, with only a slight decrease of $1\%$ in accuracy.
@@
Ablation on Models and Data. We further demonstrate the effectiveness of LoX across different aligned models. Specifically, we apply LoX to a LLaMA-2-7B model aligned with $22.5$k examples (Section 3.3) and to a Mistral-7B-v0.3 model (Jiang et al., 2023), aligned with 22.5k and 65.6k examples. We train the Mistral model for $1$ epoch instead of $3$ for LLaMA-2-7B, and keep the rest hyperparameters the same. We evaluated ASR performance on GSM8K, Alpaca, and Identity Shifting attacks, following the methodology described in Section 3.3 and Section 5.1. All models are aligned to achieve an initial ASR of 00 before fine-tuning. We use the effective rank (experiments for determining the effective rank are presented in Appendix C) and set $\alpha=1.25$ for all models, except for the Mistral model aligned with 65.6k examples, for which we use $\alpha=0.5$ due to nonsensical generated results when extrapolating further.
@@
As shown in Table 2, LoX significantly enhances the robustness of both architectures across different data sizes. For the LLaMA-2-7B model trained with 22.5k samples, the most substantial improvement is observed on the GSM8K task, where the robustness increases by $34\%$ compared to the baseline (i.e. without LoX). For the LLaMA-2-7B model trained with 65.6k samples, the greatest gain occurs under the Identity Shifting attack, with an improvement of $42\%$. For the Mistral-7B-v0.3 architecture, the largest improvements are observed in the Identity Shifting attack for both data sizes. The models aligned with 22.5k and 65.6k samples show improvements of $32.3\%$ and $33.3\%$, respectively.
@@
Ablations on Rank and Coefficient. In previous experiments, we used the definition of effective rank to determine the number of ranks to extrapolate. In this section, we examine how varying the number of extrapolated top-ranks ($6$, $100$, $500$, and the full ranking) affects robustness against the Alpaca, Identity Shifting, and GSM8K attacks. Our analysis is conducted using the LLaMA-2-7B model aligned with $65.6$k examples. In addition, we vary the extrapolation factor $\alpha$ to understand its effect on robustness. Note that extrapolating the full rank is similar to ExPO (Zheng et al., 2024).
@@
As observed in Fig. 5, low-rank approximations consistently provide the best results. For GSM8K, an ASR of $0\%$ is obtained using $k=6$ and $k=100$. For Alpaca, using $k=6$ achieves the lowest ASR of $9\%$ while consistently achieving the best performance with all extrapolation factors. These results showcase the utility of the low-rank approximations adopted by our method. For the Identity Shifting attack, while the lowest ASR of 7.3% is achieved with $k=100$, adopting low-rank approximation also achieve comparable results.
During the experiments, we observed that excessive extrapolation can destabilize the models, leading to meaningless outputs. Notably, extrapolating the full-rank matrix caused model outputs to break earlier than when extrapolating low-rank counterparts, likely due to excessive noise in the bottom-ranks. As a result, we do not report results for an extrapolation factor of $1.25$ in the full-rank setting. More details on the effects of extrapolation prior to fine-tuning can be found in Table 7 and Fig. 8, in Appendix F.
@@ 5.3 Understanding LoX via Safety Landscape
To further understand how LoX makes LLM robust, we visualize the safety landscape (Peng et al., 2024) around the aligned model, namely the ASR in a subspace of parameters. Specifically, we plot the function $F(\alpha,\beta)=\text{ASR}(\theta_{\text{align}}+\alpha d_{1}+\beta d_{2})$, where $d_{1}$ is the safety extrapolation direction, $d_{2}$ is the average of the fine-tuning directions from the aligned and extrapolated models, and $\alpha,\beta\in\mathbb{R}$. We restrict $\alpha\in[-10,10]$, $\beta\in[-8,12]$ for the Alpaca fine-tuning, and $\beta\in[-10,30]$ for the GSM8k fine-tuning, as we find that larger values lead to broken models. The projections of fine-tuned models are marked in the figure for visualization. More details on how the visualization is created are presented in Appendix G.
@@
In Fig. 6, we conduct experiments using the LLaMA-2-7B model aligned with 65.6k examples. Without LoX, the aligned model resides in a narrow valley, and therefore, the benign fine-tuning can unintendedly drag the model out of the safe region to a high ASR. After the extrapolation by LoX, the model migrates from the valley to a flatter area with similarly low ASR, away from the risky boundary. Intuitively, the flat zone implies that the model can be robustly safe upon limited perturbation. As a result, even if the same fine-tuning drags the model in the same direction, the model remains safe afterward.
@@ Appendix B Experimental details
We present the hyper-parameters for the alignment and fine-tuning applied in Section 3.3 and Section 5.1.
Alignment. We apply direct preference optimization (DPO) to safety align LLaMA-2-7B on the HH-RLHF dataset (Bai et al., 2022). We conduct multiple experiments with 22.5k, 32.8k, 45k, and 65.6k examples using the OpenRLHF codebase (Hu et al., 2024). The model is trained for $3$ epochs with a total batch size of $128$, an alignment strength $\beta=0.1$, and a learning rate of $5\times 10^{-6}$. The maximum sequence length is set to $1024$.
GSM8k Fine-tuning. We fine-tune the aligned model on GSM8K (Cobbe et al., 2021b), a dataset curated for mathematical reasoning. We use a total batch size of $20$ and train the model for $2$ epochs with a learning rate of $5\times 10^{-5}$
Alpaca Fine-tuning. Consistent with (Qi et al., 2023), we exclude safety-related examples from the training process. The model is trained for $1$ epoch with a batch size of $16$, a gradient accumulation step of $4$, a learning rate of $2\times 10^{-5}$, and a maximum gradient norm of $2$. We use a linear learning rate scheduler with a warmup step of $20$. The maximum sequence length is set to $1024$.
@@
Identity Shifting Attack. The Identity Shifting attack consists of $10$ examples designed to transform the model into an “Absolutely Obedient” Agent. We train for $10$ epochs with a batch size of $5$, a learning rate of $5\times 10^{-5}$, and a maximum gradient norm of $2$. We use a linear learning rate scheduler for training and set the maximum sequence length to $1024$. Due to the high variance observed in the ASR (see Appendix E for individual run metrics), we report the average ASR over 3 runs.
Pure Bad. The Pure Bad dataset consists of $100$ harmful examples, extracted from the Anthropic Red Teaming Dataset (Ganguli et al., 2022). We use a batch size of $5$ to train models for 5 epochs with a learning rate of $5\times 10^{-5}$.

### v3

@@ 1 Introduction
Based on this insight, we propose Low-Rank Extrapolation (LoX), a simple yet effective training-free method for enhancing the robustness of safety-aligned LLMs. LoX extrapolates the subspace of the alignment weight updates to strengthen LLMs against fine-tuning (see Fig.˜1). Through extensive experimentation, we demonstrate that our method can improve robustness against both benign and malicious fine-tuning attacks, without compromising the model’s ability to be further fine-tuned for new tasks. Furthermore, we present detailed ablation studies on the effects of the chosen subspace and the scaling coefficient. Interestingly, when inspecting the safety landscape, we observe that LoX moves the aligned model away from a narrow valley, where small perturbations can easily degrade safety, into a flatter and, therefore, safety-robust region. Our contributions are:
@@ 2 Related Work
Mitigations of Finetuning Attack. After the risks of fine-tuning were identified, several mitigation strategies were proposed (Huang et al., 2024c; Rosati et al., 2024; Wang et al., 2024; Yi et al., 2024), including (a) modifying the alignment stage to account for robustness against fine-tuning attacks (Rosati et al., 2024; Liu et al., 2024; Huang et al., 2024b; d); (b) modifying the fine-tuning stage to mitigate safety degradation (Wang et al., 2024; Zong et al., 2024; Bianchi et al., 2023; Mukhoti et al., 2023); and (c) recovering the safety from already fine-tuned models (Yi et al., 2024; Huang et al., 2024a). Another line of work focuses on improving robustness during the fine-tuning stage through techniques such as retaining base model features or applying adversarial training (Dong et al., 2021; Hou et al., 2022; Kim et al., 2023). Additional discussion regarding these methods is presented in Appendix˜A.
@@ 3.2 Low-Rank Hypothesis for Safety Model Updates
where the columns of $U$ and $V$ are the left-singular vectors and right-singular vectors, respectively. The matrix $S$ is diagonal and contains the singular values of $M$, ordered from the largest to the smallest along the main diagonal. Based on SVD, one can decompose $M$ into a summation of $r$ matrices with a rank of 111 and consequently
@@ 3.3 Fine-tuning Vanishes Safety Subspaces
Setting. We apply direct preference optimization (DPO) to safety align LLaMA-2-7B on the HH-RLHF dataset (Bai et al., 2022). We conduct multiple experiments with 22.5k, 32.8k, 45k, and 65.6k examples using the OpenRLHF codebase (Hu et al., 2024). Subsequently, we fine-tune the aligned models on GSM8K (Cobbe et al., 2021b), a dataset curated for mathematical reasoning. See Appendix˜B for hyper-parameter details.
Evaluation. In addition to our proposed metrics (i.e., $R_{\text{align}}$ and $R_{\text{ft}}$), we compute the Attack Success Rate (ASR) using the first 100100100 examples from AdvBench (Zou et al., 2023a). This allows us to quickly iterate on multiple models. Following the methodology described in Qi et al. (2023), we perform a GPT-based evaluation with a revised prompt (presented in Appendix˜H). We use multiple values of $k$ (10,100,50010,100,50010 , 100 , 500, and 200020002000) to compute the projections ${\mathrm{Proj}}_{{k}}{(.)}$. The resulting values for $R_{\text{ft}}$ and $R_{\text{align}}$ are averaged across all weight matrices.
Figure 2: Comparison of ASR difference before and after fine-tuning on GSM8K (“Safety Degradation”) versus the ratio $R_{\text{ft}}/R_{\text{align}}$. Marker size represents the number of alignment examples (22.522.522.5k, 32.832.832.8k, 454545k, 65.665.665.6k). Higher $R_{\text{ft}}/R_{\text{align}}$ correlates with lower safety degradation, suggesting the role of top-ranks in safety robustness.
[…]
Secondly, we find that in all cases $R_{\text{ft}}/R_{\text{align}}<1$ (so $R_{\text{ft}}<R_{\text{align}}$), which indicates that the proportion of safety knowledge decreases after fine-tuning. Furthermore, we observe a correlation between safety robustness and the ratio $R_{\text{ft}}/R_{\text{align}}$. For low values of $k$, large variations in $R_{\text{ft}}/R_{\text{align}}$ are correlated with larger variations in the ASR. As $k$ increases, changes in $R_{\text{ft}}/R_{\text{align}}$ become less significant, highlighting the importance of the top-ranks on the safety robustness. The exact values for the ASR, $R_{\text{ft}}$ and $R_{\text{align}}$ are presented in Appendix˜D.
@@ 4 New Method: Low-Rank Extrapolation
Based on the insights from Section˜3, we propose a new method to enhance the safety robustness of an aligned model. To maximize the flexibility and compatibility of our method, we target a general defense without making assumptions about the fine-tuning mechanisms. This new method, called Low-rank Extrapolation (LoX), strengths the safety subspaces in LLMs by extrapolating the top-ranks of $\Delta W_{align}$. More specifically, we define $W_{up}$ as
@@
where $\theta_{r}=\{W_{\text{base}}^{i}+{\mathrm{Proj}}_{{r}}{(\Delta W^{i}_{\text{align}})}\}_{i=i}^{L}$, $\text{ASR}(\theta)$ denotes the Attack Success Rate of the model of parameters $\theta$, and $\rho\in\mathbb{R}$ is a threshold which we set to 0.010.010.01. We highlight the properties of LoX:
@@ 5.1 LoX Robustifies LLMs Against Fine-tuning Attacks
To evaluate the efficacy of LoX, we fine-tune our model, aligned with 65.665.665.6k examples of the HH-RLHF dataset (Bai et al., 2022), on a variety of benign and malicious fine-tuning tasks, both before and after applying LoX. We use $k=6$ (the effective rank) and set $\alpha=1.25$. The experiment for determining the effective rank is provided in Appendix˜C.
Attack Methods. In addition to the GSM8K fine-tuning described in Section˜3.3, we also evaluate our method on Alpaca (Taori et al., 2023), Dolly (Conover et al., 2023), Identity Shifting Attack (Qi et al., 2023) and Pure Bad (Qi et al., 2023) tasks. For details about the datasets and hyper-parameters, see Appendix˜B.
@@
[…]
@@ 5.2 Ablation Study
Evaluation with Stronger Attacks. A natural question that arises is whether the obtained safety robustness can be compromised by strengthening the attack. To examine this, we conduct additional experiments on the Dolly and GSM8K datasets, using varying learning rates and training for additional epochs. For Dolly fine-tuning, we use learning rates of $2\times 10^{-5}$ and $5\times 10^{-5}$, and for GSM8k, $5\times 10^{-5}$ and $1\times 10^{-4}$. Dolly is evaluated as in Section˜5.1, while GSM8k accuracy is reported on 100 examples of the test set, for quick checkpoint assessment. Training runs for 6 epochs.
As shown in Fig.˜4, LoX improves the model’s robustness across all epochs and fine-tuning settings, with the exception of the GSM8K with $10^{-4}$ learning rate. In this case, LoX performs identically to the baseline in epochs 111 and 666, but worse in epoch 222. When inspecting the utility of the models, we observe very similar performance between LoX and the baseline, indicating that LoX does not negatively impact the model’s ability to be further adapted to new tasks.
[…]
@@
Ablation on Models and Data. We further demonstrate the effectiveness of LoX across different aligned models. Specifically, we apply LoX to a LLaMA-2-7B model aligned with 22.522.522.5k examples (Section˜3.3) and to a Mistral-7B-v0.3 model (Jiang et al., 2023), aligned with 22.5k and 65.6k examples. We train the Mistral model for 111 epoch instead of 333 for LLaMA-2-7B, and keep the rest hyperparameters the same. We evaluated ASR performance on GSM8K, Alpaca, and Identity Shifting attacks, following the methodology described in Section˜3.3 and Section˜5.1. All models are aligned to achieve an initial ASR of 0 before fine-tuning. We use the effective rank (experiments for determining the effective rank are presented in Appendix˜C) and set $\alpha=1.25$ for all models, except for the Mistral model aligned with 65.6k examples, for which we use $\alpha=0.5$ due to nonsensical generated results when extrapolating further.
@@
[…]
@@
Ablations on Rank and Coefficient. In previous experiments, we used the definition of effective rank to determine the number of ranks to extrapolate. In this section, we examine how varying the number of extrapolated top-ranks (666, 100100100, 500500500, and the full ranking) affects robustness against the Alpaca, Identity Shifting, and GSM8K attacks. Our analysis is conducted using the LLaMA-2-7B model aligned with 65.665.665.6k examples. In addition, we vary the extrapolation factor $\alpha$ to understand its effect on robustness. Note that extrapolating the full rank is similar to ExPO (Zheng et al., 2024).
@@
[…]
During the experiments, we observed that excessive extrapolation can destabilize the models, leading to meaningless outputs. Notably, extrapolating the full-rank matrix caused model outputs to break earlier than when extrapolating low-rank counterparts, likely due to excessive noise in the bottom-ranks. As a result, we do not report results for an extrapolation factor of 1.251.251.25 in the full-rank setting. More details on the effects of extrapolation prior to fine-tuning can be found in Table˜7 and Fig.˜8, in Appendix˜F.
@@ 5.3 Understanding LoX via Safety Landscape
To further understand how LoX makes LLM robust, we visualize the safety landscape (Peng et al., 2024) around the aligned model, namely the ASR in a subspace of parameters. Specifically, we plot the function $F(\alpha,\beta)=\text{ASR}(\theta_{\text{align}}+\alpha d_{1}+\beta d_{2})$, where $d_{1}$ is the safety extrapolation direction, $d_{2}$ is the average of the fine-tuning directions from the aligned and extrapolated models, and $\alpha,\beta\in\mathbb{R}$. We restrict $\alpha\in[-10,10]$, $\beta\in[-8,12]$ for the Alpaca fine-tuning, and $\beta\in[-10,30]$ for the GSM8k fine-tuning, as we find that larger values lead to broken models. The projections of fine-tuned models are marked in the figure for visualization. More details on how the visualization is created are presented in Appendix˜G.
@@
In Fig.˜6, we conduct experiments using the LLaMA-2-7B model aligned with 65.6k examples. Without LoX, the aligned model resides in a narrow valley, and therefore, the benign fine-tuning can unintendedly drag the model out of the safe region to a high ASR. After the extrapolation by LoX, the model migrates from the valley to a flatter area with similarly low ASR, away from the risky boundary. Intuitively, the flat zone implies that the model can be robustly safe upon limited perturbation. As a result, even if the same fine-tuning drags the model in the same direction, the model remains safe afterward.
@@ Appendix B Experimental details
We present the hyper-parameters for the alignment and fine-tuning applied in Section˜3.3 and Section˜5.1.
Alignment. We apply direct preference optimization (DPO) to safety align LLaMA-2-7B on the HH-RLHF dataset (Bai et al., 2022). We conduct multiple experiments with 22.5k, 32.8k, 45k, and 65.6k examples using the OpenRLHF codebase (Hu et al., 2024). The model is trained for 333 epochs with a total batch size of 128128128, an alignment strength $\beta=0.1$, and a learning rate of $5\times 10^{-6}$. The maximum sequence length is set to 102410241024.
GSM8k Fine-tuning. We fine-tune the aligned model on GSM8K (Cobbe et al., 2021b), a dataset curated for mathematical reasoning. We use a total batch size of 202020 and train the model for 222 epochs with a learning rate of $5\times 10^{-5}$
Alpaca Fine-tuning. Consistent with (Qi et al., 2023), we exclude safety-related examples from the training process. The model is trained for 111 epoch with a batch size of 161616, a gradient accumulation step of 444, a learning rate of $2\times 10^{-5}$, and a maximum gradient norm of 222. We use a linear learning rate scheduler with a warmup step of 202020. The maximum sequence length is set to 102410241024.
@@
Identity Shifting Attack. The Identity Shifting attack consists of 101010 examples designed to transform the model into an “Absolutely Obedient” Agent. We train for 101010 epochs with a batch size of 555, a learning rate of $5\times 10^{-5}$, and a maximum gradient norm of 222. We use a linear learning rate scheduler for training and set the maximum sequence length to 102410241024. Due to the high variance observed in the ASR (see Appendix˜E for individual run metrics), we report the average ASR over 3 runs.
Pure Bad. The Pure Bad dataset consists of 100100100 harmful examples, extracted from the Anthropic Red Teaming Dataset (Ganguli et al., 2022). We use a batch size of 555 to train models for 5 epochs with a learning rate of $5\times 10^{-5}$.